### v3

@@ 4 Contributions to the Literature
Our results inform monotonicity assumptions, which have become widespread. Influential work by Imbens and Angrist (1994) proposes a monotonicity assumption that assumes away compilers or defiers in the first stage of an instrumental variable model, and Manski (1997b) proposes an analogous assumption in the reduced form. Monotonicity assumptions are ubiquitous in the analysis of first stages and they are gaining traction in the analysis of reduced forms (Alsan, Cawley, Doyle, and Skelley, 2025). First stage monotonicity assumptions are useful for the interpretation of instrumental variable estimates as local average treatment effects on compliers, but they can be difficult to defend in some contexts, particularly in health care settings where interventions may be helpful for some but harmful for others. We interpret our binary outcome as “takeup” in our applications so that we can use the familiar terminology of always and never “takers” developed for the first stage (Angrist, Imbens, and Rubin, 1996), but there is nothing in our approach that requires our binary outcome to be interpretable as “takeup.” There is also nothing in our approach that requires our binary outcome to be interpretable as a “treatment” on a second stage outcome. Given concerns about side effects in medicine, important applications could involve a binary outcome that represents death.
@@
Previous approaches have revealed evidence about the joint distribution of potential outcomes that violates monotonicity, but only with additional data beyond our binary intervention and outcome. Machine learning approaches of Wager and Athey (2018) and Semenova (2024) require rich data on covariates to reveal heterogeneous effects. Analysis of side effects in medicine requires data on secondary outcomes (Bernard et al., 2001). Specification tests of instrumental variable model assumptions can reveal evidence against monotonicity in the first stage (Imbens and Rubin, 1997; Richardson and Robins, 2010; Huber and Mellace, 2012, 2015; Kitagawa, 2015; Mourifié and Wan, 2017; Machado, Shaikh, and Vytlacil, 2019), and marginal treatment effect methods can reveal evidence against monotonicity in the second stage (Björklund and Moffitt, 1987; Heckman and Vytlacil, 1999; Kowalski, 2023a, b), but such approaches require the additional structure of a two-stage model as well as data on both stages. Evidence against first stage monotonicity by Chan, Gentzkow, and Yu (2022) requires a multi-valued instrument and the ability to observe a second stage outcome for one value of the first stage outcome. We contribute to the literature on violations of monotonicity by demonstrating that, even without additional data, our design-based model of an experiment can reveal evidence about monotonicity.
@@
We view the development of our visualizations as a secondary contribution. Our visualization of potential outcomes in Figure 1 allows us to depict how the same data can arise from different distributions of potential outcomes within a Fréchet set. Our visualization of how the MLE varies with the data in Figure 3 allows us to observe patterns that would be difficult to infer from isolated empirical applications.
@@
We also offer substantive insights in actual empirical applications. In one, a payment intervention to induce pregnant women to quit smoking, we find no defiers. In another, a nudge intervention to get more people to hypothetically donate their organs, we find both compilers and the upper Fréchet bounds on defiers. In both cases, the weak evidence that we find can motivate stronger assumptions that researchers can use to learn more.
5 Implications for Future Econometric and Applied Research
@@
We see our demonstration that a design-based model can yield evidence beyond the average effect with very minimal assumptions and data as a proof-of-concept that could pave the way for the production of stronger evidence that will achieve the ultimate goal of personalizing decisions. Even if the average effect is the decision-relevant parameter in some contexts, learning about heterogeneity in the average effect is a first step toward targeting that can ultimately yield a larger average effect. In our second empirical application, under the maximum likelihood joint distribution of potential outcomes, which includes no never takers, we can deduce that the people who are untreated in intervention must be defiers, and the people who are untreated in control must be compliers. Under monotonicity, it is never possible to deduce that a given individual is a complier, and defiers are assumed away. Our hope is that the ability to look directly at some compliers and defiers under alternative assumptions motivated by our statistical decision rule will pave the way to characterize them, target interventions toward compilers and away from defiers, and improve the average effect of future interventions.
Our work suggests several areas for further econometric research. Design-based likelihoods can be derived for more complex experimental settings, including those with covariates, two-stage models, and models with alternative randomization approaches, such as matched pair, stratified, and permuted block designs. These design-based likelihoods can be used to determine exact distributions of experimental data, test statistics, or confidence intervals without resorting to simulation methods, as is common in the randomization inference literature. Approximation methods and advances in algorithmic efficiency would facilitate these calculations in large experiments. The design-based likelihood from an experiment can be integrated into a variety of choice problems beyond the specific statistical decision problem considered here. Optimal estimators under alternative utility functions or priors can be derived. Or, a design-based likelihood can be incorporated into treatment choice problems that seek to maximize average outcomes or novel targets including outcome quantiles (Guggenberger, Mehta, and Pavlov, 2024), welfare (Cui and Han, 2023), or counterfactual expected utilities (Ben-Michael, Imai, and Jiang, 2024; Christy and Kowalski, 2024b). Results from these new lines of inquiry could be used to engage with previous results on optimal experimental design that focus on estimating the average effect (Bai, 2022).
@@
Design-based methods have been around for a long time, and we have demonstrated that it is feasible to implement our design-based decision rule, but a practical impediment to widespread implementation is that applied researchers do not report enough information about their experiments to replicate them with design-based methods (Young, 2019). We use large language models to pull (openAI’s GPT-4o-mini) and categorize (openAI’s GPT-o3-mini) the randomization processes from 2,080 papers associated with randomized controlled trials from the Abdul Latif Jameel Poverty Action Lab (J-PAL). Only 61% have enough description for the large language model to confidently categorize the randomization method from the paper or associated AEA RCT Registry entry. However, of those, 78% use designs captured by design-based likelihoods that we consider. In the full set of papers, over 60% describe some mechanism through which there could be defiers. Another practical impediment to design-based methods is that instead of reporting absolute numbers, several researchers only report rates or regression results. Mandated reporting of experimental design and standard sample counts by J-PAL, medical journals, and other authorities could facilitate the application of our design-based decision rule, allowing researchers to learn more from their experiments.
@@
Appendix A  Optional Auxiliary Statistics
Figure A.1: Applications: Standard Statistics, Design-Based Maximum Likelihood Estimates, and Optional Statistics
Appendix B  Derivation of the Design-Based Likelihood
The design-based likelihood of the joint distribution of potential outcomes in the sample ${\boldsymbol{\theta}}=(\theta_{11},\theta_{10},\theta_{01},\theta_{00})$ given the data $\boldsymbol{x}$} is the probability of this realization of the data given $\boldsymbol{\theta}$}. To derive the distribution of the data, first let us define the joint distribution of potential outcomes in the intervention arm ${\boldsymbol{I}}\equiv\big(I_{11},I_{10},I_{01},I_{00}\big)$ as a random vector whose elements represent the numbers of always takers, compliers, defiers, and never takers randomized into intervention. These values are unobserved to the experimenter. However, the randomization process—the design of the experiment—implies a data generating process for the distribution of potential outcomes in the intervention arm $\boldsymbol{I}$} conditional on the distribution of potential outcomes in the sample $\boldsymbol{\theta}$}, which we can then use to derive the distribution of the data $\boldsymbol{X}$}.
@@ Appendix B  Derivation of the Design-Based Likelihood
${\mathbb{P}}\Big(I_{11}=I_{11},I_{10}=I_{10},I_{01}=I_{01},I_{00}=I_{00}\mid{\boldsymbol{\theta}}\Big)$ $=\binom{\theta_{11}}{I_{11}}\binom{\theta_{10}}{I_{10}}\binom{\theta_{01}}{I_{01}}\binom{\theta_{00}}{I_{00}}$
$\qquad\;\times p^{\sum_{j,k}i_{j,k}}(1-p)^{n-\sum_{j,k}i_{j,k}}.$ (5)
@@
${\mathbb{P}}\Big(I_{11}=I_{11},I_{10}=I_{10},I_{01}=I_{01},I_{00}=I_{00}\mid{\boldsymbol{\theta}}\Big)=\frac{\binom{\theta_{11}}{I_{11}}\binom{\theta_{10}}{I_{10}}\binom{\theta_{01}}{I_{01}}\binom{\theta_{00}}{I_{00}}}{\binom{n}{m}}$. (6)
In either case, the experimental design implies a data-generating process for the number of subjects of each type randomized into the intervention arm. While the distribution of potential outcomes in the intervention arm $\boldsymbol{I}$} is unobservable, we can use its data generating process to derive the distribution of the observable data $\boldsymbol{X}$}. Note that each subject randomized into the intervention arm with outcome $Y=1$ must be either an always taker or a complier: $X_{I1}=I_{11}+I_{10}$. Each subject randomized into the intervention arm with outcome $Y=0$ must be either a never taker or a defier: $X_{I0}=I_{00}+I_{01}$. In the control arm, those observed with outcome $Y=1$ must be either always takers that were not randomized into intervention, or defiers that were not randomized into intervention: $X_{C1}=\theta_{11}-I_{11}+\theta_{01}-I_{01}$. And finally, those in the control arm with outcome $Y=0$ must be either never takers that were not randomized into intervention, or compliers that were not randomized into intervention: $X_{C0}=\theta_{00}-I_{00}+\theta_{10}-I_{10}$. Thus, we can write the probability of the observed data $\boldsymbol{X}$} conditional on the joint distribution of potential outcomes $\boldsymbol{\theta}$} as:
@@
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}\bigg/\binom{n}{m}$ (9)
@@
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}$ (10)
This final expression is equivalent to the likelihood in Copas (1973).
Appendix C  Derivation of a Canonical Sampling-Based Likelihood
Suppose that subjects are sampled into the experiment with replacement from a population: let $q_{11}$ represent the share of the population who are always takers $q_{10}$ the share of compliers, $q_{01}$ the share of defiers , and $q_{00}$ the share never takers. Sampling with replacement implies that each subject’s vector of potential outcomes is independently and identically distributed (IID) according to these shares.
@@ Appendix C  Derivation of a Canonical Sampling-Based Likelihood
$\qquad\big(q_{11}+q_{10}\big)^{x_{I1}}\big(1-(q_{11}+q_{10})\big)^{x_{I0}}$
$\qquad\big(q_{11}+q_{01}\big)^{x_{C1}}\big(1-(q_{11}+q_{01})\big)^{x_{C0}}.$ (11)
This likelihood is equivalent to a likelihood appearing in Barnard (1947).
Alternatively, if the experiment is “completely randomized,” such that $m\leq n$ subjects are assigned to intervention by drawing names from a hat and the remainder are assigned to control, we can view the sample as two sets of Bernoulli trials. The first set of $m$ trials represent the subjects sampled into the intervention arm, where there are $X_{I1}$ treated subjects with independent probabilities of being treated equal to the probability that $Y_{I}=1$ in the population, $q_{11}+q_{10}$. The second set of nmn-m trials represent the subjects sampled into the control arm, where there are $X_{C1}$ treated subjects with independent probabilities of being treated equal to the probability that $Y_{C}=1$ in the population, $q_{11}+q_{01}$. The two sets of Bernoulli trials are independent, allowing us to write the likelihood of a completely randomized experiment as the product of two Binomial random variables $X_{I1}$ and $X_{C1}$:
@@
$\qquad\binom{n-m}{x_{C1}}\big(q_{11}+q_{01}\big)^{x_{C1}}\big(1-(q_{11}+q_{01})\big)^{n-m-x_{C1}}.$ (12)
This likelihood is equivalent to likelihoods from Barnard (1947) and Kline and Walters (2020).
In each of these likelihood functions, the joint distribution of potential outcomes in the population $\boldsymbol{q}$} appears only through the transformations $q_{11}+q_{10}$ and $q_{11}+q_{01}$, which are the marginal distributions of the potential outcomes in intervention and control in the population, respectively. Therefore, the likelihoods of any two distributions in the same Fréchet set are equal. That is, the experimental data contains no information about the joint distribution of potential outcomes in the population beyond the marginal distributions of potential outcomes.
Appendix D  Bayes Optimality of the Maximum Likelihood Decision Rule
Appendix D.1  Preliminaries on Statistical Decision Theory
@@
The decision maker chooses a decision rule that maps the observable data into (possibly) randomized guesses. [^4] We conflate here the standard definitions of “randomized decision rules” and “behavioral decision rules” (Ferguson, 1967) for expositional clarity. In settings of perfect recall, such as the setting we study here, the space of randomized and behavioral decision rules is equivalent (Kuhn, 1953). We write such a rule as $f:\boldsymbol{\mathcal{X}}\to\Delta({\boldsymbol{\Theta}})$, where $\boldsymbol{\mathcal{X}}$ is the space of possible data realizations, $\boldsymbol{\Theta}$} is the space of possible distributions of potential outcomes, and $\Delta({\boldsymbol{\Theta}})$ is the space of distributions over $\boldsymbol{\Theta}$}. Given a true distribution of potential outcomes $\boldsymbol{\theta}$}, the decision maker’s expected utility from following a decision rule $f$ is the expected value of $U(f({\boldsymbol{X}}),{\boldsymbol{\theta}})$ with respect to the experimental outcome $\boldsymbol{X}$}:
@@
Under the specified utility function, the decision maker’s expected utility is equal to the probability of guessing correctly.
@@
$EU_{\pi}(f)=\mathbb{E}_{\pi}\big[EU(f,{\boldsymbol{\theta}})\big]=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}EU(f,{\boldsymbol{\theta}})\pi({\boldsymbol{\theta}}),$
@@
Appendix D.2  Conditions for Bayes Optimality
In this appendix, we review that the result that the maximum likelihood decision rule is Bayes optimal when utility takes the form in (13) and the decision maker’s subjective prior is uniform across realizations of the distribution of potential outcomes in the sample $\boldsymbol{\theta}$}. To show this result, we first establish that the maximum a posteriori decision rule is Bayes optimal for our chosen loss function given any prior. The maximum a posteriori decision rule $f^{*}_{\pi}$ selects the maxima of the posterior distribution of $\boldsymbol{\theta}$}. Formally, let $\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})$ be the set of $\boldsymbol{\theta}$} values that maximize the posterior distribution given the observed data $\boldsymbol{x}$}, i.e.

### v4

@@ 4 Contributions to the Literature
Our results provide evidence that can support a monotonicity assumption or a specific alternative, which can be useful because monotonicity assumptions have become widespread, but they are not always realistic. Influential work by imbens1994 proposes a monotonicity assumption that assumes away compliers or defiers in the first stage of an instrumental variable model, and manski1997monotone proposes an analogous assumption in the reduced form. Monotonicity assumptions are ubiquitous in the analysis of first stages, and they are gaining traction in the analysis of reduced forms (alsan2025). First stage monotonicity assumptions are useful for the interpretation of instrumental variable estimates as local average treatment effects on compliers, and we view them as reasonable in many contexts. However, they can be difficult to defend in some contexts, particularly in health care settings that motivate our work because interventions may be helpful for some but harmful for others. We interpret our binary outcome as “takeup” in our stylized example and applications so that we can use the familiar terminology of always and never “takers” developed for the first stage (angrist1996), but the design-based likelihood is valid for other binary outcomes, like survival. Given concerns about side effects in medicine, we see potential for useful applications that have perfect compliance and a single binary outcome that represents survival. In those applications, evidence of defiers could motivate researchers to reduce the dosage of a potentially toxic intervention.
@@
Previous approaches have revealed evidence that violates monotonicity assumptions, but only with additional data beyond our binary intervention and outcome. Machine learning approaches of wager2018 and semenova2024 require data on covariates to reveal heterogeneous effects. Analysis of side effects in medicine requires data on secondary outcomes (e.g. bernard2001). Specification tests of instrumental variable model assumptions can reveal evidence against monotonicity in the first stage (imbens1997; richardson2010; huber2012; huber2015testing; kitagawa2015; mourifie2017; machado2019), and marginal treatment effect methods can reveal evidence against monotonicity in the second stage (bjorklund1987; heckman1999; kowalski2023behavior; kowalski2023reconciling), but such approaches require the additional structure of a two-stage model as well as data on both stages. Evidence against first stage monotonicity by chan2022 requires a multi-valued instrument and the ability to observe a second stage outcome for one value of the first stage outcome. We contribute to the literature on violations of monotonicity by demonstrating that, even without additional data, our design-based model of an experiment can reveal evidence for or against monotonicity.
@@
We view the development of our visualizations as a secondary contribution. Figure 1 allows us to depict how the same data can arise from different distributions of potential outcomes within a Fréchet set. Our visualization of how the MLE varies with the data in Figure 3 allows us to depict all possible results from an experiment of a given size, demonstrating that we have not engaged in selective reporting. The visualization allows us to characterize patterns that would be difficult to infer from isolated empirical applications and demonstrates the value of statistical decision theory over testing.
@@
We also offer substantive insights in real empirical applications. In one, a payment intervention to induce pregnant women to quit smoking, we find no defiers. In another, a nudge intervention to get more people to hypothetically donate their organs, we find both compliers and the upper Fréchet bounds on defiers. In both applications, the weak evidence that we find can motivate stronger assumptions that researchers can use to learn more.
5 Implications for Econometric and Applied Research
@@
We see our demonstration that a design-based model can reveal evidence beyond the average effect with very minimal assumptions and data as a proof-of-concept that could pave the way for the production of stronger evidence that will achieve the ultimate goal of personalizing decisions. Even if the average effect is the decision-relevant parameter in some contexts, learning about heterogeneity that underlies the average effect is a first step toward targeting that can ultimately yield a larger average effect. In our second empirical application, under the maximum likelihood joint distribution of potential outcomes, which includes no never takers, we can deduce that the people who are untreated in intervention must be defiers, and the people who are untreated in control must be compliers. Under monotonicity, it is never possible to deduce that a given individual is a complier, and defiers are assumed away. Our hope is that the ability to look directly at some compliers and defiers under alternative assumptions motivated by our statistical decision rule will pave the way to characterize them, target interventions toward compliers and away from noncompliers, and improve the average effect of future interventions.
Our work suggests several areas for further econometric research. Design-based likelihoods can be derived for more complex experimental settings, including those with two stages, multi-valued variables, and alternative randomization designs, such as matched pair, stratified, randomized permuted block, and designs that allow for rerandomization. A promising direction is applying design-based likelihoods to parameters traditionally studied under the lens of partial identification, such as the average treatment effect in experiments with imperfect compliance, missing data, attrition, spillovers, and measurement error. Another is using design-based likelihoods and entropy to motivate alternative parametric assumptions in auctions (jun2024) and other designed markets (abdulkadiroğlu2017). Approximation methods and advances in algorithmic efficiency could facilitate calculations in a wide array of samples. Optimal estimators under alternative priors and utility functions could be implemented with straightforward modifications. We are particularly interested in extending decision problems with “do no harm” objectives (mueller2023; benmichael2024; christy2024a; koch2025; gelman2025) to incorporate evidence from the design-based likelihood. Estimators could also be developed for alternative estimands, like potential outcome quantiles (cui2023; guggenberger2024). Finally, while previous results on optimal experimental design focus on estimating the average effect (bai2022), other designs, especially those that incorporate additional data, could reveal stronger evidence on the joint distribution of potential outcomes.
@@
Design-based methods have been around for a long time, and we have demonstrated that it is feasible to implement our design-based decision rule, but a practical impediment to widespread implementation is that applied researchers often do not report enough information about their experiments to replicate them with design-based methods (young2019). We use large language models to pull (openAI’s GPT-4o-mini) and categorize (openAI’s GPT-o3-mini) the randomization processes from 2,080 papers associated with randomized controlled trials from the Abdul Latif Jameel Poverty Action Lab (J-PAL). Only 61% have enough description for the large language model to confidently categorize the randomization method from the paper or associated AEA RCT Registry entry. However, of those, 78% use designs captured by design-based likelihoods that we consider. In the full set of papers, over 60% describe some mechanism through which there could be defiers. Another practical impediment to design-based methods is that instead of reporting absolute numbers, several researchers only report rates or regression results. Mandated reporting of experimental design and standard sample counts by J-PAL, medical journals, and other authorities could facilitate the application of our design-based decision rule, allowing researchers to learn more from their experiments.
@@
Appendix A  Derivation of the Design-Based Likelihood
The design-based likelihood of the joint distribution of potential outcomes in the sample ${\boldsymbol{\theta}}=(\theta_{11},\theta_{10},\theta_{01},\theta_{00})$ given the data $\boldsymbol{x}$} is the probability of this realization of the data given $\boldsymbol{\theta}$}. First, we define the joint distribution of potential outcomes in the intervention arm ${\boldsymbol{I}}\equiv\big(I_{11},I_{10},I_{01},I_{00}\big)$ as a random vector whose elements represent the numbers of always takers, compliers, defiers, and never takers randomized into intervention; we write its realizations as ${\boldsymbol{i}}=(i_{00},i_{10},i_{01},i_{00})$. These values are unobserved to the experimenter. However, the randomization process—the design of the experiment—implies a data generating process for the distribution of potential outcomes in the intervention arm $\boldsymbol{I}$} conditional on the distribution of potential outcomes in the sample $\boldsymbol{\theta}$}, which we can then use to derive the distribution of the data $\boldsymbol{X}$}.
@@ Appendix A  Derivation of the Design-Based Likelihood
${\mathbb{P}}\Big({\boldsymbol{I}}={\boldsymbol{i}}\mid{\boldsymbol{\theta}}\Big)=\binom{\theta_{11}}{i_{11}}\binom{\theta_{10}}{i_{10}}\binom{\theta_{01}}{i_{01}}\binom{\theta_{00}}{i_{00}}p^{\sum_{j,k}i_{j,k}}(1-p)^{n-\sum_{j,k}i_{j,k}}.$ (5)
@@
${\mathbb{P}}\Big({\boldsymbol{I}}={\boldsymbol{i}}\mid{\boldsymbol{\theta}}\Big)=\frac{\binom{\theta_{11}}{i_{11}}\binom{\theta_{10}}{i_{10}}\binom{\theta_{01}}{i_{01}}\binom{\theta_{00}}{i_{00}}}{\binom{n}{m}}$. (6)
While the distribution of potential outcomes in the intervention arm $\boldsymbol{I}$} is unobservable, we can use its data generating process to derive the distribution of the observable data $\boldsymbol{X}$}. Note that each subject randomized into the intervention arm with outcome $Y=1$ must be either an always taker or a complier: $X_{I1}=I_{11}+I_{10}$. Each subject randomized into the intervention arm with outcome $Y=0$ must be either a never taker or a defier: $X_{I0}=I_{00}+I_{01}$. In the control arm, those observed with outcome $Y=1$ must be either always takers that were not randomized into intervention, or defiers that were not randomized into intervention: $X_{C1}=\theta_{11}-I_{11}+\theta_{01}-I_{01}$. And finally, those in the control arm with outcome $Y=0$ must be either never takers that were not randomized into intervention, or compliers that were not randomized into intervention: $X_{C0}=\theta_{00}-I_{00}+\theta_{10}-I_{10}$. Thus, we can write the probability of the observed data $\boldsymbol{X}$} conditional on the joint distribution of potential outcomes $\boldsymbol{\theta}$} as:
@@
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}\bigg/\binom{n}{m}$. (9)
@@
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}$. (10)
This final expression is equivalent to the likelihood in copas1973.
Appendix B  Derivation of a Canonical Sampling-Based Likelihood
Suppose that subjects are sampled into the experiment with replacement from a population: let $q_{11}$ represent the share of the population who are always takers, $q_{10}$ the share of compliers, $q_{01}$ the share of defiers, and $q_{00}$ the share never takers. Sampling with replacement implies that each subject’s pair of potential outcomes is independently and identically distributed (IID) according to these shares.
@@ Appendix B  Derivation of a Canonical Sampling-Based Likelihood
$\qquad\times\big(q_{11}+q_{10}\big)^{x_{I1}}\big(1-(q_{11}+q_{10})\big)^{x_{I0}}$
$\qquad\times\big(q_{11}+q_{01}\big)^{x_{C1}}\big(1-(q_{11}+q_{01})\big)^{x_{C0}}.$ (11)
This likelihood is equivalent to a likelihood appearing in barnard1947.
Alternatively, if the experiment is “completely randomized,” such that $m\leq n$ subjects are assigned to intervention by drawing names from a hat and the remainder are assigned to control, we can view the sample as two sets of Bernoulli trials. The first set of $m$ trials represents the subjects sampled into the intervention arm, where there are $X_{I1}$ treated subjects with independent probabilities of being treated equal to the probability that $Y_{I}=1$ in the population, $q_{11}+q_{10}$. The second set of nmn-m trials represents the subjects sampled into the control arm, where there are $X_{C1}$ treated subjects with independent probabilities of being treated equal to the probability that $Y_{C}=1$ in the population, $q_{11}+q_{01}$. The two sets of Bernoulli trials are independent, allowing us to write the likelihood of a completely randomized experiment as the product of two Binomial random variables $X_{I1}$ and $X_{C1}$:
@@
$\qquad\times\binom{n-m}{x_{C1}}\big(q_{11}+q_{01}\big)^{x_{C1}}\big(1-(q_{11}+q_{01})\big)^{n-m-x_{C1}}.$ (12)
This likelihood is equivalent to likelihoods from barnard1947 and kline2020.
In each of these likelihood functions, the joint distribution of potential outcomes in the population $\boldsymbol{q}$} appears only through the transformations $q_{11}+q_{10}$ and $q_{11}+q_{01}$, which are the marginal distributions of the potential outcomes in intervention and control in the population, respectively. Therefore, the likelihoods of any two distributions in the same population-level Fréchet set are equal. That is, the experimental data contains no information about the joint distribution of potential outcomes in the population beyond the marginal distributions of potential outcomes.
Appendix C  Bayes Optimality of Our Proposed Rule
Appendix C.1  Preliminaries on Statistical Decision Theory
@@
The decision maker chooses a decision rule that maps the observable data into (possibly) randomized guesses. [^3] We conflate here the standard definitions of “randomized decision rules” and “behavioral decision rules” (ferguson1967) for expositional clarity. In settings of perfect recall, such as the setting we study here, the space of randomized and behavioral decision rules is equivalent (kuhn1953). We write such a rule as $f:\boldsymbol{\mathcal{X}}\to\Delta({\boldsymbol{\Theta}})$, where $\boldsymbol{\mathcal{X}}$ is the space of possible data realizations, $\boldsymbol{\Theta}$} is the space of possible distributions of potential outcomes, and $\Delta({\boldsymbol{\Theta}})$ is the space of distributions over $\boldsymbol{\Theta}$}. Given a true distribution of potential outcomes $\boldsymbol{\theta}$}, the decision maker’s expected utility from following a decision rule $f$ is the expected value of $U(f({\boldsymbol{X}}),{\boldsymbol{\theta}})$ with respect to the experimental outcome $\boldsymbol{X}$}:
@@
The last line follows from our chosen utility function, under which the decision maker’s expected utility is equal to the probability of guessing correctly.
@@
$EU_{\pi}(f)=\mathbb{E}_{\pi}\big[EU(f,{\boldsymbol{\theta}})\big]=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}EU(f,{\boldsymbol{\theta}})\pi({\boldsymbol{\theta}}),$ (14)
@@
Appendix C.2  Conditions for Bayes Optimality
In this appendix, we review the result that the maximum likelihood decision rule is Bayes optimal when utility takes the form in (13) and the decision maker’s subjective prior is uniform across realizations of the distribution of potential outcomes in the sample $\boldsymbol{\theta}$}. To show this result, we first establish that the maximum a posteriori decision rule is Bayes optimal for our chosen loss function given any prior. The maximum a posteriori decision rule $f^{*}_{\pi}$ selects the maxima of the posterior distribution of $\boldsymbol{\theta}$}. Formally, let $\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})$ be the set of $\boldsymbol{\theta}$} values that maximize the posterior distribution given the observed data $\boldsymbol{x}$}, i.e.